OpenStudy (anonymous):

Solve the following inequality and graph the solution set |x+8|>3

OpenStudy (anonymous):

\[(-\infty,11) \upsilon (-5,\infty) \]

OpenStudy (anonymous):

I think thats the answer

OpenStudy (anonymous):

you have the right idea but the wrong answer

OpenStudy (anonymous):

\[|x+8|>3\implies x+8<-3 \text{ or } x+8>3\]

OpenStudy (anonymous):

\[x+8<-3\] \[x<-11\] or \[x+8>3\] \[x>-5\] so solution is \[(-\infty, -11)\cup (-5,\infty)\]